OpenStudy (anonymous):

I need someone to check my work!! What is the sum of the arithmetic sequence 9, 14, 19, …, if there are 36 terms? answer: 3,474 What is the sum of the arithmetic sequence 308, 304, 300, …, if there are 23 terms? answer: 6,072

OpenStudy (anonymous):

For the first one, the difference between terms is 5 and there are 35 steps. 9 + 35(5) = ? For the second, the sequence decreases with each step. So answer must be less than 308, right?

OpenStudy (lgbasallote):

for the second one \[a_{23} = 308 + (23 - 1)(-4)\] \[\implies 308 -88\] \[\implies220\] \[s_n = \frac{23(308 + 220)}{2}\] \[s_n = 6,072\]

OpenStudy (anonymous):

okay ty.

OpenStudy (lgbasallote):

for the first one \[a_{36} = 9 + (36-1)(5)\] \[\implies 9 + 175\] \[\implies 184\] \[s_n = \frac{36(184 + 9)}{2}\] \[\implies 3,474\]

OpenStudy (anonymous):

ty :)